A box of chocolates contains 20 identically shaped chocolates. Four of them are filled with jelly, seven are filled with caramel, and nine are filled with nuts. What is the probability that one chocolate chosen at random is filled with caramel or nuts?
Answer by Alan3354(69443)

A box of chocolates contains 20 identically shaped chocolates. Four of them are filled with jelly, seven are filled with caramel, and nine are filled with nuts. What is the probability that one chocolate chosen at random is filled with caramel or nuts?
--------------------
Since it's "or", you add the components.
Caramel + nuts = 16
The chance is 16/20 = 4/5.
---------------------------
To look at it another way, it's the chance that it's not jelly. There are 4 jellies, so the chance of a jelly is 4/20.
20/20 - 4/20 = 16/20 = 4/5.